solve the following systems by substitution.
1. 2x+4y=10
y=3x-1

Answer by checkley77(12844)

2x+4y=10
y=3x-1 replace y with (3x-1) in the first equation & solve for x.
2x+4(3x-1)=10
2x+12x-4=10
14x=10+4
14x=14
x=14/14
x=1 ans.
y=3*1-1
y=3-1
y=2
Proof:
2*1+4*2=10
2+8=10
10=10